In Desmos: Write all transformations that occurred parent function y= square root of x

Transformed function: y= - square root x-3+1


List all transformations that occurred:

1
Expert's answer
2021-03-11T13:58:16-0500

"y=\sqrt{x}" (blue line)

"y=-\sqrt{x}" - symmetry about the x-axis (green line)

"y=-\sqrt{x-3}" - parallel translation by 3 units to the right (purple line)

"y=-\sqrt{x-3}+1" - parallel translation by 1 unit up (red line)